OpenStudy (anonymous):

How to find the value of x from 3x+15 and 2x-10

OpenStudy (anonymous):

3x + 15 = 2x - 10 First you subtract 2x from 3x. -2x -2x ----------------- x + 15 = -10 Then subtract 15 from both sides. - 15 -15 ------------ x = -25

OpenStudy (anonymous):

Ok thanks n I got another one I don't understand

OpenStudy (anonymous):

Shoot XD

OpenStudy (anonymous):

what is the sum of the interior angles of a polygon whose sides are 9,12,6 and 6

OpenStudy (jiteshmeghwal9):

subtract 2x from both sides & then add 10 both sides

OpenStudy (ajprincess):

The sum of interior angles of a polygon with four sides is always 360 despite the lengths of sides